OpenStudy (anonymous):

Choose the correct simplification of the expression -5x2(4x - 6x2 - 3). 30x4 - 20x3 + 15x2 -11x4 - x3 - 8x2 -30x4 + 20x3 - 15x2 30x4 + 20x3 + 15x2

Nnesha (nnesha):

multiply everything which is inside the parentheses by outside term and when you multiply same bases you should add their exponents

OpenStudy (triciaal):

also pay attention to the signs

OpenStudy (anonymous):

i still don't understand :(

Nnesha (nnesha):

example a(b+c) multiply b and c by outside term which is a a times b + a times c =ab+ac

Nnesha (nnesha):

multiply all 3 terms 4x,-6x^2 ,-3 by -5x^2

OpenStudy (anonymous):

Will it end up negative or positve?

Nnesha (nnesha):

you have to multiply 2 terms if one of them is negative then answer will be negative negative times negative =positive negative times positive = negative

Nnesha (nnesha):

-5x^2 times 4x =??

OpenStudy (anonymous):

so this one will end up negative because the outside term is negative?

OpenStudy (anonymous):

Cant you not combine them because one has an exponent and the other doesn't

Nnesha (nnesha):

you will get 3 terms and all of them end up with different signs bec it also depends on inside terms

Nnesha (nnesha):

yes right in order to combine them you u should have same variable with same exponents x^2 +x you can't combine these both terms bec both hve different exponents

Nnesha (nnesha):

-5x^2 times 4x = (-5 times 4)(X^2 times x) add exponents so it should be like this x^{2+1}

OpenStudy (anonymous):

I'm so confused

Nnesha (nnesha):